OpenStudy (anonymous):

Solve \[\log_{3} \][(x+5)(x-3)]=2

OpenStudy (anonymous):

In exponential form this is equivalent to solving \((x+5)(x-3) = 3^2 \implies x^2+2x-24 = 0\). Can you take things from here?

OpenStudy (anonymous):

I am supposed to come up with a solution set

OpenStudy (anonymous):

Right...so what do you get when you solve that quadratic equation?

OpenStudy (anonymous):

its \[\log_{3} \] in front of it

OpenStudy (anonymous):

Did you include the log3?

OpenStudy (anonymous):

You can rewrite logarithmic equations as exponential equations. In this case, we have that \[\log_3((x+5)(x-3)) = 2 \implies (x+5)(x-3)=3^2\implies x^2+2x-24 = 0.\]

OpenStudy (anonymous):

Oh! I understand now. So it should be: (x+6) (x-4)

OpenStudy (anonymous):

x=-6, x=4 ?

OpenStudy (anonymous):

Okay, that's good so far, so now you need to test them in the original equation to see if any of your solutions are extraneous.

OpenStudy (anonymous):

In this case, it doesn't seem like we get any extraneous solutions, so that would imply that both of the values you found are your solutions. Does this make sense? :-)

OpenStudy (anonymous):

Yes! Thank you